Shift report 20011030 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 30 2001 - 08:04:12 EST

  • Next message: Flemming Videbaek: "Follow-up on shift report (DC)"

    ------------------------------------------------------------
    Date: Tue Oct 30 08:04:12 2001, Shift: 00:00-08:00
    Supervisor: DO, with: 
    Log book#10 pages 71 - 73
    
    Shift Summary:
    ------------------------------------------------------------
    1.Machine status:
      MCR provided very good beams most of the night.
      The store started around 2:00, the ZDC coincidence
      rate was around 1000. After 4 hours, it went down to
      220. 
      At 8:00, ZDC around 150. Voted for a dump with Phenix.
    
    2.Achievements:
      Runs   Trig 6    Setting
      5574    42k      MRS 40, 6B   FS 8, 1/5 B    
      5577    44k      idem  
      5578    56k      MRS 45, 4B   FS 8, 1/5 B   
      5580    72k      MRS 45, 4B   FS 8, 1/5 A (no DC data)
      5581    13k      same setting, got back T3 and T4 (not T5)
      
    3.Detector problems:
      A problem at run start time, had to restart the DAQ
      IO and Run servers, otherwise things work really good
      until 4:30, no problems at all.
      The D3 anode currents seemed too high (58, 52, 47).
    	
      After the change in the FS polarity, started run 5579.
      3mn after, a big bad trip!!! T2, T3, T4, T5.
      I took me around 40mn to get T2, T3 and T4 back.
      Some channels refused to cooperate whatever action taken
      (turn off/on, decreased HV, turn off/on all, etc). It was
      really frustrating because of the nice store we had. 
      Finally, T5 cathode 2 kept on tripping all the time
      and T4 was unstable (cf. logbook for more details).
      I had to take the decision to run without the DCs.
    
      So, run 5580 is without DC data. Finally, could get back
      T3 and T4 to work but not T5, cathode 2 always trips.   
      Run 5581 is without T5.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 30 2001 - 08:04:59 EST